OpenStudy (anonymous):

In my book they did...1/6.28 x 100 x 2 x 10^-6 and got 796 ohms how did they get that? When I do it I get something totally different.

OpenStudy (anonymous):

You want formula ?

OpenStudy (anonymous):

anything that will show me how to get what they got

OpenStudy (anonymous):

Ok .for example R ?

OpenStudy (anonymous):

What was the problem in your book ?

OpenStudy (anonymous):

1/6.28 x 100 x 2 x 10^-6

OpenStudy (anonymous):

For example R=I.V I=V/R

OpenStudy (anonymous):

10^-6 is e .

OpenStudy (anonymous):

Xc1=1/6.28 x 100 x 2 x 10^-6=796 ohms

OpenStudy (anonymous):

What s the X?

OpenStudy (anonymous):

What is the X means ?

OpenStudy (anonymous):

reactance

OpenStudy (anonymous):

What are the data s problem ? What say every numbers in Equation ?